# PersonaMark: Personalized LLM watermarking for model protection and user attribution

### v1

@@ Abstract
The rapid development of LLMs brings both convenience and potential threats. As costumed and private LLMs are widely applied, model copyright protection has become important. Text watermarking is emerging as a promising solution to AI-generated text detection and model protection issues. However, current text watermarks have largely ignored the critical need for injecting different watermarks for different users, which could help attribute the watermark to a specific individual. In this paper, we explore the personalized text watermarking scheme for LLM copyright protection and other scenarios, ensuring accountability and traceability in content generation. Specifically, we propose a novel text watermarking method PersonaMark that utilizes sentence structure as the hidden medium for the watermark information and optimizes the sentence-level generation algorithm to minimize disruption to the model’s natural generation process. By employing a personalized hashing function to inject unique watermark signals for different users, personalized watermarked text can be obtained. Since our approach performs on sentence level instead of token probability, the text quality is highly preserved. The injection process of unique watermark signals for different users is time-efficient for a large number of users with the designed multi-user hashing function. As far as we know, we achieved personalized text watermarking for the first time through this. We conduct an extensive evaluation of four different LLMs in terms of perplexity, sentiment polarity, alignment, readability, etc. The results demonstrate that our method maintains performance with minimal perturbation to the model’s behavior, allows for unbiased insertion of watermark information, and exhibits strong watermark recognition capabilities.
@@ 1 Introduction
Since the advent of ChatGPT, large language models (LLMs) have gained significant traction, driving advancements across various fields such as automatic news composition, social media content generation, and creative writing assistance. LLMs have become indispensable tools, influencing numerous aspects of daily life [2, 10, 5, 18, 21, 40, 33, 31, 6]. Concurrently, the rise of customized and proprietary LLMs tailored for specific domains has introduced new ethical, legal, and social challenges. These include concerns around content authenticity, copyright, intellectual property (IP) protection, and the regulation of AI-generated content. Companies developing domain-specific LLMs face critical challenges in safeguarding proprietary knowledge and copyrighted material.
@@
Among the most vulnerable are personalized LLMs, which require robust copyright protections. These models, designed to replicate specific tones, personalities, or speech patterns—often aligned with the Myers-Briggs Type Indicator (MBTI) or other personal characteristics—are increasingly employed in various professional and personal contexts [29, 34, 24, 27]. As the trend towards personalized LLMs grows, ensuring the authenticity and control of their generated outputs becomes crucial. Examples of such models include Character.AI [^1] https://character.ai/, CharacterGLM[44], and ChatPLUG[30], which consistently produce outputs that reflect distinct personalities and stable character traits, including unique language styles.
As traditional post-hoc detection methods falter, text watermarking has emerged as a promising solution to protect LLM-generated content [42, 37, 7, 22]. Recent advances in text watermarking embed imperceptible signals, or watermark messages, into text data, making them detectable by algorithms but invisible to human readers[12, 43, 19, 17, 25, 41, 28, 9, 35, 38, 14, 15, 20, 7, 23]. These signals are embedded either during text generation or through post-processing. Current techniques often manipulate token distributions by dividing vocabulary into “green” and “red” lists, promoting certain tokens while suppressing others.
@@
Figure 1: Overview of our proposed PersonaMark watermark injection framework.
@@
However, these methods fall short of providing the protection and tracing capabilities found in traditional image watermarking. Personalized watermarking, where different users receive content embedded with distinct watermarks, is essential for copyright protection in customized LLMs. Yet, current techniques are limited in this regard. They typically focus on AI-generated text detection, embedding simple binary watermark signals, and fail to generate unique watermarks for different users. Even multi-bit watermarking methods suffer from limited payload capacity, which compromises robustness and text quality, ultimately restricting their ability to manage large user bases and perform effective tracing. Moreover, the reliance on token probability manipulation to inject watermarks introduces biases that undermine the training objectives of LLMs, degrading the quality of the generated text. As a result, existing methods are inadequate for achieving the same level of protection and traceability as traditional watermarking techniques, highlighting the need for more advanced approaches in the context of personalized LLMs.
@@
To address the challenges mentioned above, we propose a novel method personalized text watermarking based on sentence structure, called PersonaMark. As shown in Figure 1, PersonaMark is a plug-in framework that enables LLMs to embed unique, user-specific watermarks into generated text while maintaining high content quality and preserving model behavior. The core motivation behind PersonaMark is to overcome the limitations of existing text watermarking techniques, which often fail to provide robust user-specific watermarking without degrading text quality.
@@
PersonaMark operates by embedding personalized watermarks into text based on the structure of sentences. The process begins with the generation of $N$ candidate sentences for each prompt using a diversity beam search. Each candidate sentence’s structure is then extracted using a dependency parser and converted into a string representation. These structures are subsequently fed into a hashing function, which divides the sentence structure space into two equal parts and assigns a binary output (0 or 1) to each structure. Only sentences corresponding to a hash output of 1 are retained, with the top-ranked sentence (based on beam score) selected as the final output for the current step. This process repeats for each subsequent sentence until the entire prompt is completed. The result is a paragraph where all sentences share the “1” feature, which can be reliably detected through a statistical $Z$-test. The $Z$-test determines whether the null hypothesis ($H_{0}$) that the numbers of 0 and 1 outputs are equal should be rejected. If rejected, it indicates the presence of the watermark.
@@
To enable user-specific watermarking, the hashing function is designed to behave uniquely for each user, based on their user ID. This creates a distinct dividing plane in the sentence structure space for each user, ensuring that the watermark injection process is individualized. During detection, the dividing plane corresponding to the detected text is identified, allowing us to trace the content back to the specific user ID. This approach is highly scalable, handling large numbers of users while preserving the quality and integrity of the generated text. By focusing on sentence structure rather than token manipulation, PersonaMark maintains the naturalness of the text while providing a robust mechanism for watermarking and attribution, addressing the critical need for copyright protection in the realm of personalized LLMs.
@@
We conducted experiments on the Gemma 2B model series using three prompt sets: story writing, fake news generation, and report generation. Our method significantly outperformed existing approaches, achieving a perplexity of 2.74 compared to KGW’s 12.32 [13]. Additionally, our approach closely matched the un-watermarked model in terms of sentiment values, key information similarity, and sentence readability scores. In terms of watermark robustness, our method excelled against various attack granularities, including word- and phrase-level substitution attacks, demonstrating its effectiveness for copyright protection and user attribution. PersonaMark provides a reliable means to control the use of private LLMs and accurately identify specific users in cases of misuse or illegal activity.
We summarize our contributions as below:
@@
(1) We propose the first personalized text watermarking approach that embeds specific watermarks based on unique user IDs. This method is highly scalable, making it suitable for protecting customized and stylized LLMs with a large amount of users.
@@
(2) We utilize the sentence structures and user-specific hashing to achieve sentence-level watermarking. The watermark is closely related to language style, thus it is hard to erase when considering retaining styles. Results show that text quality is preserved and there is less influence on model behaviors, making our approach more viable for real-world applications.
(3) We introduce three indicators designed to comprehensively assess the impact of watermarking techniques on model behavior and text outputs, hoping to enhance the research on text watermarking and providing new metrics for the community.
@@ 2.3 LLM watermarking
Under the idea of adding watermarks in the process of generating text using language models, Kirchenbauer et al.[12] proposed a digital watermarking framework for large models that can embed signals that are invisible to humans but can be detected by algorithms in the generated text to mitigate the potential harm of large models. This framework achieves watermark embedding and detection by selecting a set of "green" markers and prioritizing their use during the generation process. Lee et al.[17] proposed a code generation watermarking method called SWEET, which improves the ability to detect AI generated code text by removing low entropy segments during watermark generation and detection, while significantly improving code quality preservation, surpassing all baseline methods, including post detection methods. Hu et al.[9] proposed an unbiased watermarking technique that enables the tracking and attribution of model generated text without affecting the output quality of large models (LLMs), while ensuring that users cannot detect the presence of watermarks, providing a new perspective for responsible AI development.
@@
DiPmark[35] protects data by embedding hidden information in the text generated by large models, while maintaining the original data distribution unchanged, achieving concealment, efficiency, and elasticity, making it a powerful solution for text attribution and quality protection. Yoo et al.[38] proposed a method called "Multi bit Watermark Allocation by Position" (MPAC), which embeds traceable multi bit information in the process of generating large models to address the problem of abuse in AI generated text. It effectively embeds and extracts long messages (>=32 bits) without sacrificing text quality, and has certain resistance to attacks. Wang et al.[32] proposed a novel encoding text watermarking method called Balance Marking, which is used to embed multi bit information in text generated by large models (LLMs). This method balances the probabilities of available and unavailable vocabulary by introducing a surrogate language model, effectively improving the quality and diversity of generated text without significantly increasing computational complexity. Kirchenbauer et al.[14] explored the reliability of watermarking technology in detecting and recording text generated by Large Models (LLMs) in real-world scenarios through empirical analysis. They found that watermarks can still be detected even when humans and machines rewrite the text, indicating that watermarking technology has high robustness in resisting various attacks and modifications. Liu et al.[20] proposed a semantically invariant robust watermarking method that utilizes semantic embeddings generated by auxiliary language models to determine watermark logic, achieving high-precision detection of large model generated text while achieving a good balance between attack robustness and security robustness.
@@ 4.1 Sentence structure representation
Extracting structure at the sentence level can better maintain the diversity of the text. The watermark information in this article is achieved through selective sampling of candidate sentences, without removing any words from the vocabulary, thus fully preserving the diversity at the lexical level. At the same time, the representation space of sentence structure is enormous. Even if some candidate structures are permanently removed, compared to word based approaches, the removed part of the sentences occupies a relatively small proportion of the overall potential sentence structure space. We set the structural threshold to 5, and adds watermarks to sentences with word lengths exceeding 5. If the sentence is short and the sentence structure space is small, rejecting some sentence structures may cause greater damage to the quality of the text. In terms of vocabulary size, if it is 65024 for a LLM, a segmentation of the vocabulary will make it reduce to 32512. This greatly reduces vocabulary diversity. Based on dependency syntax analysis, there are around 45 common dependency syntax labels. With a threshold of 5 when not considering the punctuation mark "punch", there are ${45^{5}}$ candidate sets in the sentence structure space based on dependency syntax analysis, with a large selectable spaces. The vast sentence structure space is much larger than the candidate word space, providing a solid foundation for imperceptible addition of watermark information and maintenance of text quality.
@@ 4.2 Sentence-level generation
In order to catch sentences as the processing objects, we improved the Beam Search generation strategy from multiple perspectives. Firstly, the ending strategy is task oriented optimized at multiple levels from the End of Sentence (EOS) perspective. Normal transformer-based LLMs are not able to stop at pre-defined punctuation during beam search process since they are trained under the special toke ’<eos>’. Our ending strategy allows the model to output only one sentence in a single computation, facilitating the addition of watermark information at the sentence level. Secondly, we use the diversity beam search to enhance the diversity of text generated content. In addition, in terms of beam scoring function, we achieve Shadow Fusion [^2] Beam search:https://opennmt.net/OpenNMT/translation/beam_searchin the beam search stage from a post-processing perspective, which can improve the effectiveness of beam search and enhance the quality of text generated. Current LLMs are usually trained on large datasets to be base models and then supervised fine-tuned to be instruction following models, thus they all often a model set. After the EOS engineering, we selectively utilize this advantage and use the base models as matching auxiliary language models to adjust the scores of candidate beams, in order to improve the text quality.
Hypothesis Filtering is used as a method to optimize the search process and is also a watermark oriented filtering task. Hypothesis filtering dynamically eliminates candidate hypotheses that are unlikely to result in high-quality output during the search process, thereby reducing computational burden and improving search efficiency. This method is particularly suitable for language models with a large number of possible outputs, and is also particularly suitable for the framework that needs to generate a set of candidate sentences in this paper. The filtering rules include multiple dimensions: 1. length limitation, which only retains assumptions with lengths within a certain range. This helps to avoid outputs that are longer or shorter than expected. 2. Repetition punishment: if a hypothesis contains words that have already appeared in previous steps, its score can be reduced or filtered out directly. 3. Prior knowledge. We filter out assumptions that violate common sense or rules based on specific domain knowledge, which is implemented by a separate checker. The rule checker requires a certain degree of flexibility, so it is independently completed at the end of the processing steps. After Hypothesis Filtering, the obtained candidate sentence set will be encoded and filtered as watermark addition objects.
@@
4.3 Watermark-driven hashing and personalization
Figure 2: Overview of the personalized watermark detection process.

### v2

@@ Abstract
The rapid advancement of customized Large Language Models (LLMs) offers considerable convenience. However, it also intensifies concerns regarding the protection of copyright/confidential information. With the extensive adoption of private LLMs, safeguarding model copyright and ensuring data privacy have become critical. Text watermarking has emerged as a viable solution for detecting AI-generated content and protecting models. However, existing methods fall short in providing individualized watermarks for each user, a critical feature for enhancing accountability and traceability. In this paper, we introduce PersonaMark, a novel personalized text watermarking scheme designed to protect LLMs’ copyrights and bolstering accountability. PersonaMark leverages sentence structure as a subtle carrier of watermark information and optimizes the generation process to maintain the natural output of the model. By employing a personalized hashing function, unique watermarks are embedded for each user, enabling high-quality text generation without compromising the model’s performance. This approach is both time-efficient and scalable, capable of handling large numbers of users through a multi-user hashing mechanism. To the best of our knowledge, this is a pioneer study to explore personalized watermarking in LLMs. We conduct extensive evaluations across four LLMs, analyzing various metrics such as perplexity, sentiment, alignment, and readability. The results validate that PersonaMark preserves text quality, ensures unbiased watermark insertion, and offers robust watermark detection capabilities, all while maintaining the model’s behavior with minimal disruption.
@@ 1 Introduction
Since the inception of ChatGPT, Large Language Models (LLMs) have garnered significant attention, spearheading advancements across various domains, including automatic news composition, social media content generation, and creative writing assistance. These models have become essential tools, impacting numerous facets of daily life [5, 14, 9, 24]. In parallel, the growth of customized and proprietary LLMs tailored to specific fields has introduced new ethical, legal, and social challenges. These challenges encompass concerns surrounding content authenticity, copyright, intellectual property (IP) protection, and the regulation of AI-generated content. Organizations developing domain-specific LLMs encounter substantial hurdles in protecting proprietary knowledge and copyrighted material.
@@
Among the most susceptible are personalized Large Language Models (LLMs), which necessitate robust copyright protections. These models are designed to emulate specific tones, personalities, or speech patterns—often aligned with the Myers-Briggs Type Indicator (MBTI) or other personal characteristics—and are increasingly utilized in various professional and personal contexts [32, 35, 26, 1, 29]. As the trend towards personalized LLMs grows, ensuring the authenticity and control of their generated outputs becomes crucial. Examples of such models include Character.AI [1], CharacterGLM[45], and ChatPLUG[33], which consistently produce outputs that reflect distinct personalities and stable character traits, including unique language styles.
@@
Figure 1: Overview of our proposed PersonaMark watermark injection framework. The italicized and underlined sentences are watermarked sentences.
Customized models, when commercialized and made accessible to the public, present significant concerns regarding potential misuse. For example, a personalized LLM trained to act as a famous host could provide users with interesting interactive experience, but it can also be used to campaign for someone, such as helping Trump win votes. Additionally, these models could potentially be misused to generate pornographic content, which is unacceptable. Whether pertaining to hosts, artists, or politicians, there is a collective intention to prevent models trained with personal data from being used to produce misinformation and mislead the public. Furthermore, there exists a risk of copyrighted information being misappropriated and misused. For example, a pharmaceutical company specializing in drugs for mental illnesses may use proprietary data to train drug discovery models for research and development purposes. However, these models might also be co-opted for the illegal production of substances, such as novel psychoactive drugs [4, 30, 2]. In scenarios of leakage and misuse, it is imperative that companies possess the capability to trace such activities to specific users or employees and hold individuals accountable.
As traditional post-hoc detection methods falter, text watermarking has emerged as a promising solution to protect LLM-generated content [43, 38, 11]. Recent advances in text watermarking embed imperceptible signals, or watermark messages, into text data, making them detectable by algorithms but invisible to human readers[16, 44, 21, 20, 27, 42, 31, 13, 36, 39, 17, 18, 22, 11, 25]. These signals are embedded either during text generation or through post-processing. Current techniques often manipulate token distributions by dividing vocabulary into “green” and “red” lists, promoting certain tokens while suppressing others.
@@
However, these methods lack the robust protection and tracing capabilities characteristic of traditional image watermarking. Personalized watermarking, which involves embedding distinct watermarks into content for different users, is crucial for copyright protection in customized LLMs. Yet, current works are insufficient to meet this needs. They predominantly focus on detecting AI-generated text and rely on embedding simple binary watermark signals, failing to produce unique watermarks for individual users. Even multi-bit watermarking techniques suffer from limited payload capacity[23], which compromises both robustness and text quality, thereby restricting their ability to manage extensive user bases[40] and to conduct effective tracing. Moreover, the manipulation of token probabilities to inject watermarks introduces biases that conflict with the training objectives of LLMs, ultimately degrading the quality of the generated text. Consequently, existing methods are inadequate for achieving the same level of protection and traceability as traditional watermarking techniques, highlighting the need for more advanced approaches in the context of personalized LLMs.
@@
To address the challenges mentioned above, we propose a novel method personalized text watermarking based on sentence structure, termed PersonaMark. As illustrated in Figure 1, PersonaMark is a plug-in framework that enables LLMs to embed unique, user-specific watermarks into generated text while maintaining high content quality and preserving model behavior. The core motivation behind PersonaMark is to overcome the limitations of existing text watermarking techniques, which frequently fall short in providing robust user-specific watermarking without compromising text quality.
@@
PersonaMark operates by embedding personalized watermarks into text based on the structure of sentences. The process begins with the generation of $N$ candidate sentences for each prompt using a diverse beam search. The structure of each candidate sentence is extracted using a dependency parser and is converted into a string representation. These structures are then input into a hashing function, which partitions the sentence structure space into two equal parts, assigning a binary output (0 or 1) to each. Only sentences corresponding to a hash output of 1 are retained, with the top-ranked sentence (based on beam score) selected as the final output for the current step. This process repeats for each subsequent sentence until the entire prompt is completed. The result is a paragraph where all sentences share the “1” feature, which can be reliably detected through a statistical $Z$-test. The $Z$-test determines whether the null hypothesis ($H_{0}$) that the numbers of 0 and 1 outputs are equal, should be rejected. If rejected, it indicates the presence of the watermark.
@@
To enable user-specific watermarking, the hashing function is designed to behave uniquely for each user, based on their user ID. This customization creates a distinct dividing plane in the sentence structure space for each user, ensuring that the watermarking process is individualized. During detection, the dividing plane corresponding to the detected text is identified, which allows tracing the content back to the specific user ID. This approach is notably scalable, capable of managing a large number of users while maintaining the quality and integrity of the generated text. By focusing on sentence structure rather than token manipulation, PersonaMark maintains the naturalness of the text while providing a robust mechanism for watermarking and attribution, addressing the critical need for copyright protection in the realm of personalized LLMs.
@@
We conducted experiments across four model series using three prompt datasets: story writing, fake news generation, and report generation. Our method significantly outperformed existing approaches, achieving a perplexity of 2.74 compared to the baseline KGW’s 12.32 [16]. Furthermore, our approach closely matched the un-watermarked model in terms of sentiment values, key information similarity, and sentence readability scores, perserving the original model behaviors. In terms of watermark robustness, our method excelled against various attack granularities, demonstrating its effectiveness for copyright protection and user attribution. Notably, under 40% rate of word-level substitution attack, the AUC score of watermark detection remains 0.97. PersonaMark provides a reliable means to control the use of private LLMs and accurately identify specific users in cases of misuse or illegal activity.
Contributions: We summarize our contributions as below:
@@
(1) We propose the personalized text watermarking approach that embeds specific watermarks based on unique user IDs. This method is highly scalable, making it suitable for protecting customized and stylized LLMs with a large amount of users with fast attribution speed.
@@
(2) We utilize the sentence structures and user-specific hashing to achieve sentence-level watermarking. The watermark is intricately linked to language style, making it difficult to remove without altering inherent styles. Results indicate that text quality is well-preserved with minimal impact on model behaviors, thereby enhancing the approach’s viability for real-world applications.
(3) We introduce three indicators designed to comprehensively assess the impact of watermarking techniques on model behavior and text outputs. Extensive experiments results demonstrate the superiority of our proposed method in terms of perplexity, sentiment polarity, key information consistency, readability and detection accuracy.
@@ 2.3 LLM watermarking
Kirchenbauer et al.[16] proposed a digital watermarking framework for large models that can embed signals that are invisible to humans but can be detected by algorithms in the generated text to mitigate the potential harm of large models. This framework(referred to as KGW in the following sections) achieves watermark embedding and detection by selecting a set of "green" markers and prioritizing their use during the generation process. Lee et al.[20] proposed a code generation watermarking method called SWEET, which improves the ability to detect AI generated code text by removing low entropy segments during watermark generation and detection, while significantly improving code quality preservation, surpassing all baseline methods, including post detection methods. Hu et al.[13] proposed an unbiased watermarking technique that enables the tracking of model generated text without affecting the output quality of large models (LLMs), while ensuring that users cannot detect the presence of watermarks, providing a new perspective for responsible AI development.
@@
DiPmark[36] protects data by embedding hidden information in the text generated by large models, while maintaining the original data distribution unchanged, achieving concealment, efficiency, and elasticity, making it a powerful solution for text attribution and quality protection. Yoo et al.[39] proposed a method called "Multi bit Watermark Allocation by Position" (MPAC), which embeds traceable multi bit information in the process of generating large models to address the problem of abuse in AI generated text. It effectively embeds and extracts long messages (32 bits). Although it holds potential applicability in user attribution scenarios, its detection accuracy diminishes substantially with an increase in bit number, as the vocabulary is partitioned into more segments rather than the original two-part division employed by KGW[16]. User attribution necessitates the management of a substantial number of accounts while ensuring that robustness and detection accuracy remain uncompromised as the user base expands. These are precisely the features of our method, as our initial design was specifically tailored for multi-user environments. Wang et al.[34] proposed a novel encoding text watermarking method called Balance Marking, which is used to embed multi bit information in text generated by large models (LLMs). This method balances the probabilities of available and unavailable vocabulary by introducing a surrogate language model, effectively improving the quality and diversity of generated text without significantly increasing computational complexity. Kirchenbauer et al.[17] explored the reliability of watermarking technology in detecting and recording text generated by Large Models (LLMs) in real-world scenarios through empirical analysis. They found that watermarks can still be detected even when humans and machines rewrite the text, indicating that watermarking technology has high robustness in resisting various attacks and modifications. Liu et al.[22] proposed a semantically invariant robust watermarking method that utilizes semantic embeddings generated by auxiliary language models to determine watermark logic, achieving high-precision detection of large model generated text while achieving a good balance between attack robustness and security robustness.
@@ 4 Method
Figure 2: Overview of the personalized watermark detection process.
@@ 4.1 Sentence structure representation
Extracting structure at the sentence level can better maintain the diversity of the text. The watermark information in this article is achieved through selective sampling of candidate sentences, without removing any words from the vocabulary, thus fully preserving the diversity at the lexical level. At the same time, the representation space of sentence structure is enormous. Even if some candidate structures are permanently removed, compared to word based approaches, the removed part of the sentences occupies a relatively small proportion of the overall potential sentence structure space. We set the structural threshold to 5, and adds watermarks to sentences with word lengths exceeding 5. If the sentence is short and the sentence structure space is small, rejecting some sentence structures may cause greater damage to the quality of the text. In terms of vocabulary size, if it is 65024 for a LLM, a segmentation of the vocabulary will make it reduce to 32512. This greatly reduces vocabulary diversity. Taking a mainstream dependency parsing system, spaCy [^1] https://spacy.io/ as example, there are around 45 common dependency syntax labels based on dependency syntax analysis(‘en_core_web_sm’). With a threshold of 5 when not considering the punctuation mark "punch", there are ${45^{5}}$ candidate sets in the sentence structure space based on dependency syntax analysis. Although the actual sentence structure possibilities are constrained by linguistic rules and semantic plausibility, the range of selectable options remains substantial when considering Martin Cutts’ recommendation that the average sentence length should ideally fall between 15 and 20 words[8]. The vast sentence structure space is much larger than the candidate word space, providing a solid foundation for imperceptible addition of watermark information and maintenance of text quality.
@@ 4.2 Sentence-level generation
In order to catch sentences as the processing objects, we improved the Beam Search generation strategy from multiple perspectives. Firstly, the ending strategy is task oriented optimized at multiple levels from the End of Sentence (EOS) perspective. Normal transformer-based LLMs are not able to stop at pre-defined punctuation during beam search process since they are trained under the special token ‘eos’. Our ending strategy allows the model to output only one sentence in a single computation, facilitating the addition of watermark information at the sentence level. Secondly, we use the diversity beam search to enhance the diversity of text generated content. In addition, in terms of beam scoring function, we achieve Shadow Fusion [^2] Beam search: https://opennmt.net/OpenNMT/translation/beam_search in the beam search stage from a post-processing perspective, which can improve the effectiveness of beam search and enhance the quality of text generated. Current LLMs are usually trained on large datasets to be base models and then supervised fine-tuned to be instruction following models, thus they all often a model set. After the EOS engineering, we selectively utilize this advantage and use the base models as matching auxiliary language models to adjust the scores of candidate beams, in order to improve the text quality.
To improve quality of output and to optimize the search process, hypothesis filtering is implemented, which dynamically eliminates candidate hypotheses that are unlikely to result in high-quality outputs. This method is particularly suitable for language models with a large number of possible outputs, as well as the framework that needs to generate a set of candidate sentences in this paper. The filtering rules include three dimensions: L, R, and P. L: Length limitation, which helps to avoid outputs that are longer or shorter than expected. R: Repetition punishment. If a hypothesis contains words that have already appeared in previous steps, its score can be reduced or filtered out directly. P: Prior knowledge. We filter out assumptions that violate common sense or pre-defined rules based on specific domain knowledge, which is implemented by a separate checker. The rule checker is independently completed at the end of the processing steps to achieve flexibility for potential changes. After Hypothesis Filtering, the obtained candidate sentence set will be encoded and filtered as watermark addition objects.
@@
4.3 Signature-driven hashing and personalization